# Auto-Calibration and Biconvex Compressive Sensing with Applications to Parallel MRI

### v1

@@ Abstract
We study an auto-calibration problem in which a transform-sparse signal is compressive-sensed by multiple sensors in parallel with unknown sensing parameters. The problem has an important application in pMRI reconstruction, where explicit coil calibrations are often difficult and costly to achieve in practice, but nevertheless a fundamental requirement for high-precision reconstructions. Most auto-calibrated strategies result in reconstruction that corresponds to solving a challenging biconvex optimization problem. We transform the auto-calibrated parallel sensing as a convex optimization problem using the idea of ‘lifting’. By exploiting sparsity structures in the signal and the redundancy introduced by multiple sensors, we solve a mixed-norm minimization problem to recover the underlying signal and the sensing parameters simultaneously. Robust and stable recovery guarantees are derived in the presence of noise and sparsity deficiencies in the signals. For the pMRI application, our method provides a theoretically guaranteed approach to self-calibrated parallel imaging to accelerate MRI acquisitions under appropriate assumptions. Developments in MRI are discussed, and numerical simulations using the analytical phantom and simulated coil sensitives are presented to support our theoretical results.
Keywords: Self-calibration, Compressive sensing, Convex optimization, Random matrices, MRI.
@@ 1 Introduction
We frequently encounter challenges with imperfect sensing, where accurately calibrated sensors are critical for high-precision measurements or reconstructions. For many applications, explicit calibration is often difficult and expensive to carry out in practice, posing a major roadblock in many scientific and technological endeavors where accurate and reliable measurements are essential. The idea of auto-calibration (calibration-less/self-calibration) avoids the difficulties and inaccuracy associated with explicit estimations by equipping the sensors and systems with the capability to automatically derive the sensing information from its collected signals while performing the intended system function. Auto-calibration is a field of research that finds successful applications across diverse fields such as wireless communication, computer vision, remote sensing, biomedical imaging and more. This field encompasses a wide range of techniques, including blind deconvolution, blind phase retrieval, blind rotation estimation, and direction-of-arrival estimation, among others.
The inverse problem for auto-calibration is generally difficult to solve. In practice, most self-calibration algorithms rely on joint estimation of the signals of interest and the calibration parameters using standard techniques such as maximum likelihood estimation, alternating minimization or convex/nonconvex optimization. In this work, we extend the line of work that combines convex optimization and compression sensing in solving various auto-calibration tasks [36][11][4][20]. A common scheme in such approaches includes first transforming the biconvex constraints to convex constraints using the idea of ‘lifting’ [11]; then, drawing on ideas from compressive sensing to exploit redundancy/sparsity priors in the signal model and/or the calibration model while being consistent with the convex constraints.
In this work, we consider a type of auto-calibration problem where the same underlying signal is being sensed by multiple sensors in parallel with unknown calibration parameters. Specifically, we are concerned with the following problem:
@@
where $\mathbf{y}_{i}\in\mathbb{C}^{L\times 1}$ are the measurements, $\mathbf{F}_{\Omega}\in\mathbb{C}^{L\times N}\text{ with }L\leq N$ is the partial Fourier matrix with rows belongs to a subset $\Omega\in[N]$, $\mathbf{s}_{i}\in\mathbb{C}^{N\times 1}$ is the vector of unknown calibration parameters, $\mathbf{x}\in\mathbb{C}^{N\times 1}$ is the our signal of interest, $\boldsymbol{\omega}_{i}\in\mathbb{C}^{L\times 1}$ is the additive noise. We have $C$ sets of measurements. Here $\odot$ is the point-wise multiplication. Our goal is to simultaneously reconstruct the signal x and sensor profiles (calibration parameters) $\textbf{h}_{i}$. Such a forward model has important applications e.g. in parallel MRI reconstruction where one collects partial Fourier (k-space) measurements of an underlying image using multiple receiver coils where each one has nonuniform responses.
Before proceeding with further analyzing the model, we briefly discuss the identifiability and dimension of the problem. First, if $\mathbf{s}_{i}$ and $\mathbf{x}$ form a pair of solutions to (1), then for any complex vector $\mathbf{v}$ with nonzero entries, $\mathbf{s}_{i}\odot\mathbf{v}$ and $\mathbf{x}\odot 1/\mathbf{v}$ is a pair of solutions. Furthermore, with $L\leq N$, (1) has $CL$ number of measurements and $(C+1)N$ number of unknowns. Thus, it is impossible to recover all $\mathbf{s}_{i}$ and $\mathbf{x}$ without making further assumptions.
Of particular interest is the case when $\mathbf{x}$ is transform-sparse, a characteristic often observed in real-world signals [^1] This notion can be easily generalized to approximately transform-sparse.. That is, there is some known sparsifying transformation $\boldsymbol{\Psi}\in\mathbb{C}^{N\times N}$ such that $\boldsymbol{\Psi}\textbf{x}$ is n-sparse with $n<N$. Common choices for $\boldsymbol{\Psi}$ include the well-known DCT basis and the wavelet basis. Despite the sparsity assumption, the available measurements are still fewer than the unknowns (i.e., $CL\leq CN+n$). We impose an additional subspace assumption, $\mathbf{s}_{i}=\mathbf{B}\mathbf{h}_{i}$, which can be interpreted as that the calibration parameters all lie in some known subspace spanned by a tall matrix $\mathbf{B}$. Practical choices for such $\mathbf{B}$ can be low-order polynomial basis or low-frequency sinusoidal basis to promote smoothness in the coil profiles. Altogether, the forward model becomes:
@@
Here, $\mathbf{B}\in\mathbb{C}^{N\times k},k<N$, $\mathbf{\Psi}\in\mathbb{C}^{N\times N}$ are known.
The forward model holds significant relevance in blind pMRI reconstructions. Additionally, although the above model may appear straightforward and specific, providing a rigorous analysis on it proves to be non-trivial. Our goal is to understand the requirements on the number of samples $L$ and the number of parallel measurements $C$ for a robust and stable recovery of all $\mathbf{h}_{i}$ and $\mathbf{x}$. The primary challenge for analysis comes from the inherent structure of the forward matrix, which requires us to make reasonable and careful assumptions regarding randomness that align with the specific application in mind.
@@ 1.1 Notation and outline
We denote vectors and matrices by bold font letters (e.g. x or X) and scalars by regular font or Greek symbols. For any vector $\textbf{x}\in\mathbb{C}^{N}$ and set $S\subset[N]$, define $\textbf{x}_{S}=\mathcal{P}_{S}\textbf{x}\in\mathbb{C}^{N}$ the orthogonal projection of x onto set $S$. (e.g. $\textbf{x}_{S}(i)=\textbf{x}(i),i\in S$ and $\textbf{x}_{S}(i)=0,\text{for }i\in S^{c}$). Similarly, define the orthogonal projection of matrix $\textbf{X}\in\mathbb{C}^{N_{1}\times N_{2}}$ to set $T\subset[N_{1}]\times[N_{2}]$ by $\textbf{X}_{T}=\mathcal{P}_{T}\textbf{X}\in\mathbb{C}^{N_{1}\times N_{2}}$, such that $\textbf{X}_{T}(i,j)=\textbf{X}(i,j),(i,j)\in T$ and $\textbf{X}_{T}(i,j)=0,\text{otherwise}$. We denote the N by N identity matrix as $\textbf{I}_{N}$. For any complex vector x or matrix X, $\bar{\textbf{x}}$ or $\bar{\textbf{X}}$ means their conjugate respectively. $vec(\textbf{X})$ means the vectorization of a matrix X in the column-wise order into a vector. We denote the circular convolution as *** and the kronecker product as $\otimes$. For any set $S$, we denote $S^{c}$ as its complementary set and $|S|=\text{Card}(S)$. When we refer to a matrix as generic, we imply that the entries within that matrix/vector are independently sampled from a standard Gaussian distribution. Informally, we refer to a matrix $X\in\mathbb{C}^{N_{1}\times N_{2}}$ being incoherent if the 2-norm of its rows does not vary too much.
The paper is organized as following: In section 2 we briefly discuss our contributions in relation to the state of the art in self-calibration and compressive sensing MRI. In section 3 we introduce the problem setup. We present our main results in section 4 and provide the details of the proofs in section 6. Numerical experiments are shown in section 5.
@@ 2 Related work
Our work is grounded in a real-world application concerning auto-calibrated pMRI. We draw general ideas from compressive sensing, convex optimization, self-calibation and applications in MRI reconstruction. The work is inspired by papers related to bringing self-calibration problems into the framework of biconvex compressive sensing, such as PhaseLift, SparseLift [36][20], and by blind deconvolution using lifting by [4].
We first introduce the concept of ‘lifting’, which serves as an important tool to in our proposed method. Consider any bi-linear measurements of two unknown signals $\mathbf{x}$ and $\mathbf{z}$ of the form
$y_{i}=<\mathbf{x},\mathbf{b_{i}}><\mathbf{z},\mathbf{\bar{a}_{i}}>,i=1\cdots m,$ (3)
where $\mathbf{x},\mathbf{b}_{i}\in\mathbb{C}^{N_{1}\times 1},\mathbf{z},\mathbf{a}_{i}\in\mathbb{C}^{N_{2}\times 1}$. Define matrix $\mathbf{X}:=\mathbf{x}\mathbf{z}^{T}$, we can write the bilinear equations as a linear equation with respect to $\mathbf{X}$ using some linear algebra,
$<\mathbf{x},\mathbf{b_{i}}><\mathbf{z},\mathbf{\bar{a}_{i}}>=\mathbf{b}_{i}^{*}\mathbf{x}\mathbf{z}^{T}\mathbf{a}_{i}=\mathbf{b}_{i}^{*}\mathbf{X}\mathbf{a}_{i}=(\mathbf{a}_{i}^{T}\otimes\mathbf{b}_{i}^{*})\operatorname{vec}(\mathbf{X}).$ (4)
Hence, by defining linear operator $\mathcal{A}:\mathbb{C}^{N_{1}\times N_{2}}\rightarrow\mathbb{C}^{m}$, the systems of equations becomes
@@
The original bilinear equations of dimension $(N_{1}+N_{2})$ is ‘lifted’ to linear equations of underlying dimension $N_{1}\times N_{2}$.
2.1 Related work in auto-calibration and compressive sensing
There exist two primary approaches to address the bilinear optimization problems: one involves convexifying it through methods such as linearization or lifting [4][36][35][20][11]. The other involves optimization in the natural parameters space using alternating minimization and gradient descent algorithms[30][33][46]. Various auto-calibration problems of similar setups have been explored, mainly differing from each other in the following two assumptions:
  • The assumptions made about the signal and calibration parameters: whether they are sparse individually or jointly, sparse in the natural parameter space or after transformations, subspace constrains, or restricted by sign constraints.
  • The selection of the forward model: whether it employs generic sensing matrices (e.g., Gaussian) or structured sensing matrices implied by the specific application cases.
We briefly survey several works that have influenced our work, employing a combination of ’lifting’, sparsity priors, and subspace assumptions. The SparseLift framework [36] is concerned with another type of bi-linear self-calibration problem pertaining sparse signals. The authors consider recovery from the following bi-convex measurements:

### v2

@@ Abstract
We study an auto-calibration problem in which a transform-sparse signal is acquired via compressive sensing by multiple sensors in parallel, but with unknown calibration parameters of the sensors. This inverse problem has an important application in pMRI reconstruction, where the calibration parameters of the receiver coils are often difficult and costly to obtain explicitly, but nonetheless are a fundamental requirement for high-precision reconstructions. Most auto-calibration strategies for this problem involve solving a challenging biconvex optimization problem, which lacks reconstruction guarantees. In this work, we transform the auto-calibrated parallel compressive sensing problem to a convex optimization problem using the idea of ‘lifting’. By exploiting sparsity structures in the signal and the redundancy introduced by multiple sensors, we solve a mixed-norm minimization problem to recover the underlying signal and the sensing parameters simultaneously. Our method provides robust and stable recovery guarantees that take into account the presence of noise and sparsity deficiencies in the signals. As such, it offers a theoretically guaranteed approach to auto-calibrated parallel imaging in MRI under appropriate assumptions. Applications in compressive sensing pMRI are discussed, and numerical experiments using real and simulated MRI data are presented to support our theoretical results.
Keywords: Self-calibration, Compressive sensing, Convex optimization, Random matrices, Parallel MRI.
@@ 1 Introduction
We frequently encounter challenges with imperfect sensing, where accurately calibrated sensors are critical for high-precision measurements and reconstructions. For many applications, explicit calibration is often difficult and expensive to carry out in practice, posing a major roadblock in many scientific and technological endeavors where accurate and reliable measurements are essential. The idea of auto-calibration (calibration-less/self-calibration) avoids the difficulties and inaccuracy associated with explicit estimations by equipping the sensors and systems with the capability to automatically derive the sensing information from its collected signals while performing the intended system function. Auto-calibration is a field of research that finds successful applications across diverse fields such as wireless communication, computer vision, remote sensing, biomedical imaging, and more. This field encompasses a wide range of techniques, including blind deconvolution, blind phase retrieval, blind rotation estimation, and direction-of-arrival estimation, among others.
The inverse problem for auto-calibration is generally difficult to solve. In practice, most self-calibration algorithms rely on joint estimation of both the signals of interest and the calibration parameters using standard techniques such as maximum likelihood estimation, alternating minimization, or convex/nonconvex optimization. In this work, we extend the line of work that combines convex optimization and compression sensing in solving various auto-calibration tasks [ling_strohmer_2015, Candès_Eldar_Strohmer_Voroninski_2013, Ahmed_Recht_Romberg_2014, flinth_2017]. A common scheme in such approaches includes first transforming the biconvex constraints to convex constraints using the concept of ‘lifting’ [Candès_Eldar_Strohmer_Voroninski_2013]; then, drawing on ideas from compressive sensing to exploit redundancy/sparsity priors in the signal model and/or the calibration model while being consistent with the convex constraints.
In this work, we consider a type of auto-calibration problem where the same underlying signal is sensed by multiple sensors in parallel with unknown calibration parameters. Specifically, we are concerned with the following problem:
@@
where $\mathbf{y}_{i}\in\mathbb{C}^{L\times 1}$ are the measurements, $\mathbf{F}_{\Omega}\in\mathbb{C}^{L\times N}\text{ with }L\leq N$ is the partial Fourier matrix with rows belonging to a subset $\Omega\subset[N]$, $\mathbf{s}_{i}\in\mathbb{C}^{N\times 1}$ is the vector of unknown calibration parameters, $\mathbf{x}\in\mathbb{C}^{N\times 1}$ is our signal of interest, $\boldsymbol{\omega}_{i}\in\mathbb{C}^{L\times 1}$ is the additive noise. We have $C$ sets of measurements. Here, $\odot$ is the point-wise multiplication. Our goal is to simultaneously reconstruct the signal x and the parameters related to the sensor profiles (calibration parameters) $\textbf{h}_{i}$.
Such a forward model appears, for instance, in blind parallel MRI (pMRI). Specifically, the $i$-th coil measurement $\mathbf{y}_{i}$ is a noisy sample of the spatial Fourier transform of an object $\mathbf{x}$, weighted by an unknown coil sensitivity $\mathbf{s}_{i}$ that corresponds to the $i$-th sensing coil. In parallel imaging, $C$ coils are used to take measurements simultaneously for the same underlying object $\mathbf{x}$. The joint estimation of coil sensitivities and the object of interest leads to the nonlinear inverse problem of the form (1).
Before proceeding with further analyzing the model, we briefly discuss the identifiability and dimension of the problem. First, if $\mathbf{s}_{i}$ and $\mathbf{x}$ form a pair of solutions to (1), then for any complex vector $\mathbf{v}$ with nonzero entries, $\mathbf{s}_{i}\odot\mathbf{v}$ and $\mathbf{x}\odot 1/\mathbf{v}$ are a pair of solutions. Furthermore, with $L\leq N$, (1) has $CL$ number of measurements and $(C+1)N$ number of unknowns. Thus, it is impossible to recover all $\mathbf{s}_{i}$ and $\mathbf{x}$ without making further assumptions.
Of particular interest is the case when $\mathbf{x}$ is transform-sparse, a characteristic often observed in real-world signals  [^1] This notion can be easily generalized to approximately transform-sparse.. That is, there is some known sparsifying transformation $\boldsymbol{\Psi}\in\mathbb{C}^{N\times N}$ such that $\boldsymbol{\Psi}\textbf{x}$ is n-sparse with $n<N$. Common choices for $\boldsymbol{\Psi}$ include the well-known DCT basis and the wavelet basis. Despite the sparsity assumption, the available measurements are still fewer than the number of unknowns (i.e., $CL\leq CN+n$). We impose an additional subspace assumption, $\mathbf{s}_{i}=\mathbf{B}\mathbf{h}_{i}$, which can be interpreted as that the calibration parameters all lie in some known subspace spanned by a tall matrix $\mathbf{B}$. Practical choices for such $\mathbf{B}$ can be a low-order polynomial basis or a low-frequency sinusoidal basis. Such choices promote smoothness in the sensing parameters. Altogether, the forward model becomes:
@@
Here, $\mathbf{B}\in\mathbb{C}^{N\times k},k<N$, $\mathbf{\Psi}\in\mathbb{C}^{N\times N}$ are known transformations.
Although the above model may appear straightforward and specific, providing a rigorous analysis is nontrivial. The primary challenge arises from carefully analyzing the incoherence of the forward matrix, which involves the interactions between $\mathbf{F}_{\Omega}$, $\mathbf{B}$, and $\boldsymbol{\Psi}$. Our goal is to understand the requirements on the number of samples $L$ and the number of channels $C$ for robust and stable recovery of all $\mathbf{s}_{i}$ and $\mathbf{x}$. We will discuss particular choices of $\mathbf{F}_{\Omega}$, $\mathbf{B}$, and $\boldsymbol{\Psi}$ that align with the pMRI application in detail in Section  2.
An important observation is that Parallel Imaging (PI) studies can sometimes appear “noisy”. This is largely because the primary objective of PI techniques is to reduce imaging time, which leads to acquiring fewer data points per coil due to the need to reduce imaging time. Consequently, this reduction results in a lower signal-to-noise ratio (SNR). As a result, the SNR of PI sequences is typically lower than that of comparable non-PI sequences [Aja-Fernández_Vegas-Sánchez-Ferrero_Tristán-Vega_2014, Robson_Grant_Madhuranthakam_Lattanzi_Sodickson_McKenzie_2008].
The SNR of PI sequences is affected by the same factors that influence non-PI sequences, such as field strength, magnet hardware, coil loading, the type of tissue being imaged, pulse sequence, timing parameters, voxel volume, the number of phase-encoding steps, receiver bandwidth, and the number of signals averaged. Additionally, the SNR for PI sequences ($\text{SNR}_{\text{parallel}}$) is further reduced by two extra factors, as shown in the equation below:
$\text{SNR}_{\text{parallel}}=\text{SNR}_{\text{non-PI}}/(\sqrt{R}\cdot g),$
where $R$ denotes the reduction or acceleration factor for undersampling, and $g$ represents a spatially dependent term known as the geometry factor. The term $g$ relates to the number, size, and orientation of the surface coil elements and can be considered a measure of coil separation [Aja-Fernández_Vegas-Sánchez-Ferrero_Tristán-Vega_2014, Robson_Grant_Madhuranthakam_Lattanzi_Sodickson_McKenzie_2008].
@@ 1.1 Notation and outline
We denote vectors and matrices by bold font letters (e.g. x or X) and scalars by regular font or Greek symbols. We denote the circular convolution as * and the Kronecker product as $\otimes$. For any integer $N$, we denote the set $\{1,\cdots,N\}$ as $[N]$. For any vector $\textbf{x}\in\mathbb{C}^{N}$ and set $S\subset[N]$, define $\textbf{x}_{S}=\mathcal{P}_{S}\textbf{x}\in\mathbb{C}^{N}$ the orthogonal projection of x onto set $S$. (e.g. $\textbf{x}_{S}(i)=\textbf{x}(i),i\in S$ and $\textbf{x}_{S}(i)=0,\text{for }i\in S^{c}$). Similarly, define the orthogonal projection of matrix $\textbf{X}\in\mathbb{C}^{N_{1}\times N_{2}}$ to set $T\subset[N_{1}]\times[N_{2}]$ by $\textbf{X}_{T}=\mathcal{P}_{T}\textbf{X}\in\mathbb{C}^{N_{1}\times N_{2}}$, such that $\textbf{X}_{T}(i,j)=\textbf{X}(i,j),(i,j)\in T$ and $\textbf{X}_{T}(i,j)=0,\text{otherwise}$. We denote the N by N identity matrix as $\textbf{I}_{N}$. For any complex vector x or matrix X, $\bar{\textbf{x}}$ or $\bar{\textbf{X}}$ means their conjugate, respectively. $vec(\textbf{X})$ means the vectorization of a matrix X in the column-wise order into a vector. For any set $S$, we denote $S^{c}$ as its complementary set and $|S|=\text{Card}(S)$. Informally, we refer to a matrix $X\in\mathbb{C}^{N_{1}\times N_{2}}$ being incoherent if the 2-norm of its rows does not vary too much.
The paper is organized as follows: In Section 2, we briefly discuss our contributions in relation to the state of the art in auto-calibration and compressive sensing pMRI. In Section 3, we introduce the problem setup. We present our main results in Section 4 and provide the details of the proofs in Section 6. Numerical experiments are shown in Section 5.
@@ 2 Related work
The work is inspired by papers related to bringing auto-calibration problems into the framework of biconvex compressive sensing. Many auto-calibration problems can be seen as jointly recovering the signal and the calibration parameters from bilinear measurements. Specifically, they can be seen as variations of recovering two unknown signals $\mathbf{x}$ and $\mathbf{z}$ from bilinear measurements,
$y_{i}=\langle\mathbf{x},\mathbf{b_{i}}\rangle\langle\mathbf{z},\mathbf{\bar{a}_{i}}\rangle,i=1\cdots m,$ (3)
where $\mathbf{x},\mathbf{b}_{i}\in\mathbb{C}^{N_{1}\times 1},\mathbf{z},\mathbf{a}_{i}\in\mathbb{C}^{N_{2}\times 1}$.
Various auto-calibration problems of similar setups have been explored, including blind deconvolution in [Ahmed_Recht_Romberg_2014], direction-of-arrival estimation as seen in [ling_strohmer_2015], blind deconvolution and demixing [flinth_2017], phase retrieval via matrix completion [Candès_Eldar_Strohmer_Voroninski_2013], among others. The analysis of the auto-calibration problems generally differs in the following two aspects:
  • Additional assumptions must be made about the signal, the calibration parameters, or both to address the ill-posed nature of the problem. Popular assumptions may vary in several ways, either regarding the signal, the sensing parameters, or both: whether they assume sparsity individually or jointly, sparsity in the natural parameter space or after transformations, subspace constraints, or restrictions imposed by sign constraints.
  • Specific application cases impose different forward models due to physical or other constraints. For example, the choice of sensing matrices - whether sub-Gaussian (e.g., $b_{i}$ or $a_{i}$ are Gaussian), structured random (e.g., taking random samples from the Fourier transform) or deterministic.
Two primary approaches to solving the bilinear optimization problem are as follows: one involves convexifying it through methods such as linearization or lifting [Ahmed_Recht_Romberg_2014, ling_strohmer_2015, Ling_Strohmer_2018, flinth_2017, Candès_Eldar_Strohmer_Voroninski_2013]. The other involves optimization in the natural parameter space using alternating minimization and gradient descent algorithms [Lee_Tian_Romberg_2018, Li_Ling_Strohmer_Wei_2019, Qu_Li_Zhu_2019].
For a general discussion on the issue of injectivity and the principle of identifiability for bi-linear compressive sensing, one may refer to [Kech_Krahmer_2017, Choudhary_Mitra_2013].
2.1 Related work in auto-calibration and compressive sensing
It seems natural and tempting to recover $(\textbf{h}_{1},\cdots,\textbf{h}_{C},\textbf{z})$ obeying (2) by solving the optimization problem
$\min\limits_{\textbf{h}_{1},\cdots,\textbf{h}_{C},\textbf{z}}\sum\limits_{i}\|\textbf{y}_{i}-\textbf{F}_{\Omega}(\textbf{B}\textbf{h}_{i}\odot\boldsymbol{\Psi}^{*}\textbf{z})\|_{2}^{2}.$ (4)
However, the above nonlinear least squares is a challenging optimization problem since it is highly nonconvex and most of the available algorithms, such as alternating minimization and gradient descent, may suffer from getting easily trapped in some local minima.
A popular convex optimization approach involves transforming the bi-convex problem (3) to the convex problem using the idea of ‘lifting’. Define the matrix $\mathbf{X}:=\mathbf{x}\mathbf{z}^{T}$, we can write the bilinear equations as a linear equation with respect to $\mathbf{X}$ using some linear algebra,
$\langle\mathbf{x},\mathbf{b_{i}}\rangle\langle\mathbf{z},\mathbf{\bar{a}_{i}}\rangle=\mathbf{b}_{i}^{*}\mathbf{x}\mathbf{z}^{T}\mathbf{a}_{i}=\mathbf{b}_{i}^{*}\mathbf{X}\mathbf{a}_{i}=(\mathbf{a}_{i}^{T}\otimes\mathbf{b}_{i}^{*})\operatorname{vec}(\mathbf{X}).$ (5)
Hence, by defining linear operator $\mathcal{A}:\mathbb{C}^{N_{1}\times N_{2}}\rightarrow\mathbb{C}^{m}$, the system of equations becomes
@@ 2.1 Related work in auto-calibration and compressive sensing
The original bilinear equations of dimension $(N_{1}+N_{2})$ are ‘lifted’ to linear equations of underlying dimension $N_{1}\times N_{2}$.
An important auto-calibration problem pertaining to sparse signals stems from the task of image reconstruction using randomly coded masks [Bahmani_Romberg_2015, Tang_Recht_2014]. The forward model for this auto-calibration problem is given by:
@@
where $\mathbf{h}\in\mathbb{C}^{k}$, $\mathbf{y}\in\mathbb{C}^{L}$, $\mathbf{A}\in\mathbb{C}^{L\times N},L\leq N$ and $\|\mathbf{x}\|_{0}=n$. In this model, the matrix A acts as the sensing matrix. The differences between the forward model (7) and the current work (1) are as follows: 1. The blind sensing (i.e., $\text{diag}(\mathbf{Bh})$) is applied after compressed sensed signals (i.e., $\mathbf{Ax}$). 2. The signal is sparse in its natural domain rather than being transform-sparse.